The school of business tells you the WACC for Realty Assist Inc. is 14%, with the before-tax cost of debt of 10.77%. The school also tells you the required return on equity is 18%.

You've found that the tax rate is 21%, then what proportion of a firm is equity financed?

70.26%

54.00%

77.78%

57.86%
Answer #1

Weight of Equity = We

Weight of Debt = Wd

We + Wd = 1

Wd = 1 - WE

WACC = 14%

befor-tax cost of debt = kd = 10.77%

tax rate = t = 21%

Cost of Equity = ke = 18%

WACC is calculated using the formula:

WACC = We*ke + Wd*kd*(1-t)

14% = We*18% + (1-We)*10.77%*(1-21%)

14% = We*18% + (1-We)*8.5083%

14% = We*18% + 8.5083% - We*8.5083%

14% - 8.5083% = We*(18% - 8.5083%)

5.4917% = We*9.4917%

We = 5.4917%/9.4917% = 57.8579179704374% ~ 57.86%

Proportion of equity = We = 57.86%

Answer -> 57.86%